### v1

# PUCP-Metrix: A Comprehensive Open-Source Repository of Linguistic Metrics for Spanish

Abstract
Linguistic features remain essential for interpretability and tasks involving style, structure, and readability, but existing Spanish tools offer limited coverage. We present PUCP-Metrix, an open-source repository of 182 linguistic metrics spanning lexical diversity, syntactic and semantic complexity, cohesion, psycholinguistics, and readability. PUCP-Metrix enables fine-grained, interpretable text analysis. We evaluate its usefulness on Automated Readability Assessment and Machine-Generated Text Detection, showing competitive performance compared to an existing repository and strong neural baselines. PUCP-Metrix offers a comprehensive, extensible resource for Spanish, supporting diverse NLP applications.
PUCP-Metrix: A Comprehensive Open-Source Repository of Linguistic Metrics for Spanish
[…]
1 Introduction
Linguistic features have gained renewed importance in explainable NLP, particularly for tasks requiring interpretability, stylistic sensitivity, or attention to surface-level properties. Despite advances in end-to-end neural models, recent work shows that handcrafted or derived features remain essential in applications such as AI-generated text detection (Kumarage et al., 2023; Ciccarelli et al., 2024; Petukhova et al., 2024), educational NLP (Mizumoto and Eguchi, 2023; Hou et al., 2025; Atkinson and Palma, 2025), and readability assessment (Zeng et al., 2024; Liu et al., 2025). In automated essay scoring, for instance, models incorporating linguistic features offer more transparent and pedagogically meaningful evaluations (Hou et al., 2025). These trends highlight the need for robust, modular repositories of linguistic metrics that can complement deep models.
Beyond NLP applications, these repositories also support linguistic research, offering standardized, quantifiable descriptions of texts across genres, registers, and proficiency levels (Jiang, 2016; Kuiken, 2023). They enable empirical analyses of morphosyntactic variation, cohesion, or lexical sophistication, and facilitate cross-linguistic comparisons.
Existing tools have demonstrated the value of this approach. For instance, Coh-Metrix (McNamara et al., 2010) provides extensive metrics for English across various linguistic levels. Similar resources include NILC-Metrix for Portuguese (Leal et al., 2023), Coh-Metrix-Esp for Spanish (Quispesaravia et al., 2016), and MultiAzterTest for Spanish, English, and Basque (Bengoetxea and Gonzalez-Dios, 2021).
In this work, we introduce PUCP-Metrix, a new open-source toolkit for extracting linguistic metrics from Spanish texts. It expands the range of available metrics across lexical, syntactic, discourse, psycholinguistic, and readability dimensions. In addition, we demonstrate its utility in two downstream tasks: Automated Readability Assessment and Machine-Generated Text Detection.
Our main contributions are:
  • PUCP-Metrix, a comprehensive and extensible open-source repository of linguistic metrics for Spanish, featuring metrics not available in existing resources. [^1] The code is available at https://github.com/iapucp/pucp-metrix.
  • An empirical study evaluating its usefulness in Automated Readability Assessment and Machine-Generated Text Detection.
2 Related Work
Linguistic analysis tools have played a key role in understanding and quantifying text complexity. Coh-Metrix McNamara et al. (2010), a widely used tool for English, provides metrics capturing lexical, syntactic, semantic, and discourse characteristics of texts. These metrics support applications from educational assessment to psycholinguistic research, offering a detailed view of text complexity. Inspired by this framework, similar tools have been developed for other languages, adapting metrics to reflect language-specific features.
For Portuguese, NILC-Metrix Leal et al. (2023) provides a comprehensive set of over 200 metrics covering lexical, syntactic, semantic, discourse, and psycholinguistic dimensions, enabling detailed text analysis for educational and linguistic research. In Spanish, both Coh-Metrix-Esp Quispesaravia et al. (2016) and MultiAzterTest Bengoetxea and Gonzalez-Dios (2021) offer comparable capabilities. Coh-Metrix-Esp adapts the original Coh-Metrix to Spanish, implementing 45 linguistic features and is applied in a Automated Readability Assesment task. MultiAzterTest combines over 125 linguistic and stylistic features with machine learning classifiers to evaluate text complexity in Spanish, English, and Basque.
3 PUCP-Metrix
3.1 System Design
We used an open-source implementation of Coh-Metrix for the Spanish language Quispesaravia et al. (2016) as a starting point for our work. To implement new metrics, we analyzed the metrics in the tools described in Section 2 and consulted their implementation details when available.
3.2 Linguistic Categories and Metrics
For our processing needs, including tokenization, dependency parsing, and POS tagging, we utilized the NLP library spaCy. We developed custom pipelines to extract linguistic metrics from the texts efficiently. Following these steps, we compiled a collection of 182 linguistic metrics for Spanish texts. The complete list is available at Appendix A.
  • Descriptives: 27 indicators that capture general statistics of the text, such as number of words, number of sentences, number of paragraphs, minimum and maximum length of sentences, average word length.
  • Lexical Diversity: 22 indicators measure the diversity of the text’s vocabulary, including the type-token ratio for various word categories (nouns, verbs, etc.), noun density, verb density, adverb density, adjective density, the Maas Index Mass (1972), MLTD, and vocd McCarthy Philip M (2010). Our implementation extends these measures with type-token ratios for additional word categories and their lemmatized forms. Key indicators include the following:
    • MTLD (Measure of Textual Lexical Diversity): Addresses TTR’s length sensitivity by calculating the average length of sequential word segments that maintain a certain TTR threshold, providing more stable measures across varying text lengths McCarthy Philip M (2010).
    • VOCd (Vocabulary Complexity Diversity): Estimates vocabulary richness through curve-fitting techniques on random samples, offering insights into the probability of encountering new word types McCarthy Philip M (2010).
    • Maas Index: A logarithmic transformation that provides an alternative measure of lexical diversity, particularly useful for comparing texts of different lengths Mass (1972).
  • Readability: 7 indicators that represent how difficult to understand the text is, such as Flesch Grade Level, Brunet Index, Gunning Fog Index, Honore’s Statistic, SMOG Grade, The Szigriszt-Pazos Perspicuity Index and Readability µ.Among the important measures are:
    • Flesch Grade Level (Fernández-Huertas adaptation): Adapted for Spanish texts, this measure estimates the grade level required for comprehension.
    • Szigriszt-Pazos Perspicuity Index: A Spanish-specific readability measure that evaluates text clarity, offering insights into Spanish text comprehensibility.
    • SMOG Grade: Estimates the years of education required to understand a text by analyzing polysyllabic words (3+ syllables).
    • Gunning Fog Index: Calculates readability by considering both sentence length and complex word percentage, estimating the education level needed for comprehension.
    • Honore’s Statistic: Measures vocabulary richness by analyzing hapax legomena (words appearing only once).
    • Readability µ: A statistical measure that evaluates text complexity through letter distribution patterns.
  • Syntactic Complexity: 12 indicators, reflecting the structural intricacy of text, such as proportion of sentences with 1-7 clauses, minimal edit distances of words, POS tags and lemmas. Following Coh-Metrix, our implementation extends minimal edit distance measures to POS tags and lemmatized forms, providing comprehensive syntactic variation analysis.
  • Psycholinguistics: 30 indicators, these reflect psycholinguistic properties of words, specifically how they are understood by humans: concreteness, imageability, familiarity, age of acquisition, valence and arousal. These psycholinguistic properties were collected from the EsPal database Duchon et al. (2013) and works from Stadthagen-Gonzalez et al. (2017):
    • Concreteness: Measures the degree to which words refer to tangible, physical objects versus abstract concepts. Higher concreteness values indicate words that are easier to visualize and process cognitively.
    • Imageability: Assesses how easily words can evoke mental images. Words with higher imageability are processed more quickly and remembered more easily.
    • Familiarity: Evaluates how well-known words are to speakers. Familiar words are processed faster and require less cognitive effort.
    • Age of Acquisition: Measures the age at which words are typically learned. Earlier acquired words are processed more automatically and efficiently.
    • Valence: Assesses the emotional positivity or negativity of words. Valence influences emotional processing and memory formation.
    • Arousal: Measures the emotional intensity or activation level of words. Arousal affects attention and memory consolidation.
  • Word Information: 24 indicators, more detailed word-level statistics, such as: number of nouns, number of verbs, number of adverbs, number of adjectives and number of content words.
  • Referential Cohesion: 12 indicators, serve to measure the interconnections within a text: noun overlap, argument overlap, stem overlap, content word overlap and anaphor overlap.
  • Textual Simplicity: 4 indicators, measure the simplicity of the text using the ratio of short or large sentences, such as: proportion of short sentences, proportion of medium sentences, proportion of long sentences, proportion of very long sentences.
  • Semantic Cohesion: 8 indicators, assessing the degree of semantic relatedness between different parts of the text, such as: LSA overlap of adjacent sentences, LSA overlap of all sentences, LSA overlap of adjacent paragraphs.
  • Word Frequency: 16 indicators, various measurements involving the Zipf’s frequency for different kinds of words, such as rare nouns count, rare verbs count, rare adverbs count, rare content words count and mean word frequency.
  • Syntactic Pattern Density: 14 indicators, reflecting the density of various syntactic elements, such as: noun phrase density, verb phrase density, negative expressions density, coordinating conjunctions density and subordinating conjunction density.
  • Connectives: 6 indicators, measuring the use of words or phrases that establish logical, temporal, or other relationships between different parts of the text, such as: casual connectives incidence, logical connectives incidence, adversative connectives incidence, temporal connectives incidence, additive connectives incidence, all connectives incidence.
3.3 Comparison with Existing Tools
Table 1 shows the number of linguistic metrics implemented in Coh-Metrix-Esp, MultiAzterTest and PUCP-Metrix (ours). PUCP-Metrix provides a broader coverage of linguistic metrics compared to CohMetrix-Esp and MultiAzterTest, comprising a total of 182 metrics across 13 categories. Notably, PUCP-Metrix includes metrics in categories that are entirely missing or underrepresented in the other tools, such as Semantic Cohesion, Textual Simplicity, and Psycholinguistics, with 8, 4, and 30 metrics, respectively. This way, PUCP-Metrix can capture higher-level discourse, cognitive readability, and psycholinguistic properties.
Furthermore, PUCP-Metrix distributes its metrics more evenly across lexical, syntactic, semantic, and psycholinguistic dimensions. This comprehensive and balanced coverage allows for a more detailed and nuanced characterization of texts, making PUCP-Metrix better suited for in-depth linguistic analysis and a wide range of NLP applications.
Category CohMetrix-Esp MultiAzterTest PUCP-Metrix (ours)
Descriptive 11 22 27
Referential Cohesion 12 10 12
Lexical Diversity 2 20 22
Readability 1 1 7
Connectives 6 12 6
Syntactic Complexity 2 19 12
Pattern Density 3 0 14
Semantic Cohesion 0 0 8
Word Information 11 32 24
Word Frequency 0 15 16
Textual Simplicity 0 0 4
Psycholinguistics 0 0 30
Word Semantic Information 0 4 0
Semantic Overlap 0 6 0
Total 48 141 182
Table 1: Number of linguistic metrics per Category for each tool.
4 Applications
In order to verify the usefulness of PUCP-Metrix, we use it for two tasks where linguistic metrics have proven to be helpful in past work. In particular, we select Automated Readability Assessment (ARA) and Machine-Generated Text Detection.
4.1 Automated Readability Assessment (ARA)
We follow an approach similar to that of Vásquez-Rodríguez et al. (2022). The original work introduced a benchmark for ARA of Spanish texts. The authors combined multiple corpora labeled by language proficiency levels and proposed 2-label and 3-label classification schemes.
In contrast, our study comprises four publicly available datasets —CAES, Coh-Metrix-Esp, Kwiziq, and HablaCultura— to ensure reproducibility and open accessibility. We adopt the same label mappings described in the paper, adapting all texts to two readability classification schemas: 2-label (simple, complex) and 3-label (basic, intermediate, advanced). The dataset’s descriptions and the labeling strategy can be found in Appendix B.
Overall, the dataset contains 32,167 instances, distributed across the four sources as follows: 31,149 from CAES, 100 from Coh-Metrix-Esp, 206 from Kwiziq, and 713 from HablaCultura.
We experiment with two readability classification schemas mentioned before. All experiments are performed at the document level [^2] We use the same texts that come from the available resources. The corpus is divided into 80% training, 10% validation, and 10% test sets, stratified by label. We evaluate models using Precision, Recall, and F1-score.
We implement a baseline model based on the Spanish variant of RoBERTa (named RoBERTa-BNE) Fandiño et al. (2022) [^3] The model was available at https://huggingface.co/PlanTL-GOB-ES/roberta-base-bne. We fine-tune RoBERTa-BNE for both the 2-label and 3-label classification tasks using the aforementioned splits.
4.2 Machine-Generated Text Detection
We adopt the AuTexTification 2023 shared task dataset Sarvazyan et al. (2023), which comprises over 160,000 texts in English and Spanish across five domains: tweets, reviews, news, legal, and how-to articles and generated by both human and large language models (machine).
For our experiments, we focus on the Machine-generated Text Detection task, which consists of identifying if a text has been created by a human or a machine. The task includes 26,996 human-generated instances and 25,195 machine-generated instances, totaling 52,191 instances. More details about the dataset can be found in Appendix B.
Following the shared task settings of AuTexTification and in line with the ARA task, we adopt RoBERTa-BNE as our baseline. It is fine-tuned on the official training splits and evaluated on the corresponding test splits to ensure comparability.
For both tasks, we trained various machine learning models: Logistic Regression (LR), XGBoost (XGB), Support Vector Machines (SVM) and Random Forest (RF) on the metrics extracted with both MultiAzter and PUCP-Metrix.
5 Results and Discussion
5.1 Automated Readability Assesment
Table 2 compares PUCP-Metrix, MultiAzter, and RoBERTa-BNE on two-label ARA/complexity classification. PUCP-Metrix slightly outperforms MultiAzter across simple and complex texts, achieving an overall F1 of 97.46 with XGBoost versus 97.24 for MultiAzter. In addition, XGBoost consistently yields the highest F1 scores, with Random Forest also performing competitively, while Logistic Regression and SVM score slightly lower.
RoBERTa-BNE achieves the best overall F1 of 98.30, indicating that although PUCP-Metrix captures rich linguistic cues, deep contextual models excel at detecting subtle semantic patterns.
Model Simple Complex F1
P R F1 P R F1
Multiazter
LR 96.42 97.27 96.85 91.75 89.37 90.54 93.70
XGB 98.05 99.20 98.62 97.57 94.19 95.85 97.24
SVM 96.51 97.32 96.91 91.89 89.62 90.74 93.82
RF 97.25 99.29 98.26 97.76 91.72 94.64 96.45
PUCP-Metrix
LR 96.68 97.65 97.16 92.87 90.11 91.47 94.31
XGB 98.38 99.08 98.73 97.22 95.18 96.19 97.46
SVM 96.60 97.69 97.14 92.97 89.86 91.39 94.27
RF 97.45 99.20 98.32 97.52 92.34 94.86 96.59
RoBERTa-BNE 99.04 99.24 99.14 97.76 97.16 97.46 98.30
Table 2: Results on 2-label ARA/Complexity Classification task
Table 3 compares PUCP-Metrix, MultiAzter, and RoBERTa-BNE on the 3-label ARA task. PUCP-Metrix again performs slightly better than Multiazter, reaching an overall F1 of 96.72 with XGBoost versus 96.56 for Multiazter, being XGBoost the model that achieves the highest scores.
Similarly to previous results, PUCP-Metrix does not surpass RoBERTa-BNE; that achieves the best overall F1 of 98.13, with near-perfect performance on Basic and Intermediate texts and strong results on Advanced ones.
Model Basic Intermediate Advanced F1
P R F1 P R F1 P R F1
Multiazter
LR 91.43 92.56 91.99 85.66 86.30 85.97 83.00 74.20 78.36 85.44
XGB 97.62 98.59 98.10 96.43 96.59 96.51 98.48 91.87 95.06 96.56
SVM 90.54 93.08 91.79 85.29 85.71 85.50 84.72 68.55 75.78 84.36
RF 96.32 98.07 97.18 94.38 94.93 94.66 98.37 85.16 91.29 94.38
PUCP-Metrix
LR 92.25 92.85 92.55 86.35 86.71 86.53 82.02 77.39 79.64 86.24
XGB 97.68 98.59 98.13 97.16 96.59 96.88 96.72 93.64 95.15 96.72
SVM 91.10 93.55 92.31 86.06 85.63 85.85 82.72 71.02 76.43 84.86
RF 95.55 98.18 96.85 95.11 93.77 94.44 97.63 87.28 92.16 94.48
RoBERTa-BNE 99.30 99.24 99.27 98.83 98.42 98.63 95.50 97.53 96.50 98.13
Table 3: Results on 3-label ARA/Complexity Classification task
5.2 Machine-Generated Text Detection
Table 4 shows the performance of various machine learning models using the metrics provided by PUCP-Metrix, and Multiazter. Also, it shows the performance of a RoBERTa-BNE model fine-tuned on the AuTexTification dataset and the overall performance of RoBERTa-BNE and the best model reported at the shared task.
In general, PUCP-Metrix consistently outperforms MultiAzter across classifiers. For human texts, PUCP-Metrix increases F1 scores from 42–51 (Multiazter) to 60–66, and for machine texts from 70–73 to 71–76, showing its ability to capture linguistic and structural cues critical for detecting human-written content. Tree-based models, particularly XGBoost and Random Forest, leverage PUCP-Metrix most effectively, achieving the highest overall F1 scores.
Compared to RoBERTa-BNE, PUCP-Metrix achieves more balanced performance across classes. While RoBERTa-BNE attains very high precision for human texts (93.96), its recall is low (37.86), yielding an F1 below PUCP-Metrix’s best result. This indicates that contextual embeddings may miss the diversity of human writing, whereas interpretable linguistic metrics maintain robust detection across both classes.
Furthermore, PUCP-Metrix slightly surpasses the best model reported in the shared task (F1 70.77), suggesting that integrating linguistic features with neural models could further improve classification performance.
Finally, we conduct an analysis about what are the linguistic metrics that are more important for classification. In general, Machine-generated text detection relies on features related to frequency, readability, and cohesion, while ARA tasks prioritize descriptive, syntactic, and simplicity features. Details are provided in Appendix C.
Model Human Machine F1
P R F1 P R F1
Multiazter
LR 70.52 30.28 42.37 61.84 89.93 73.29 57.83
XGB 68.10 39.73 50.18 63.98 85.19 73.08 61.63
SVM 70.43 30.74 42.80 61.95 89.73 73.30 58.05
RF 62.08 43.98 51.49 63.82 78.62 70.45 60.97
PUCP-Metrix
LR 71.09 55.93 62.61 70.02 81.90 75.49 69.05
XGB 71.34 61.36 65.97 72.33 80.38 76.14 71.06
SVM 71.04 56.05 62.66 70.06 81.82 75.48 69.07
RF 63.57 58.24 60.79 68.85 73.44 71.07 65.93
RoBERTa-BNE 93.96 37.86 53.97 66.48 98.06 79.24 66.61
RoBERTa-Autex* - - - - - - 68.52
Best model* - - - - - - 70.77
Table 4: Results on AuTexTification. *The authors of the shared task only provide F1 in the report.
6 Tool Usage
PUCP-Metrix can be installed via pip:
pip install iapucp-metrix
To use the library, we need to import the Analyzer class and call compute_metrics to compute all metrics. The function supports multiprocessing through spaCy, allowing us to specify the number of workers and the batch size.
from iapucp_metrix.analyzer import Analyzeranalyzer = Analyzer()texts = ["Este es mi ejemplo."]metrics_list = analyzer.compute_metrics( texts, workers=4, batch_size=2)for i, metrics in enumerate(metrics_list): print(Readability (Fernández-Huertas):) print(f"{metrics[’RDFHGL’]:.2f}")
The output of the code described above is:
Readability (Fernández-Huertas):201.86
In addition, PUCP-Metrix supports computing metrics grouped by linguistic categories (via compute_grouped_metrics), enabling users to analyze model behavior across dimensions such as lexical, syntactic, and semantic features.
7 Conclusion and Future Work
PUCP-Metrix provides a linguistically rich set of 182 metrics for Spanish, offering broader coverage and a larger metric set than previous resources. Empirical evaluations demonstrate its effectiveness in ARA and Machine-generated text detection tasks. Models trained on these metrics match or outperform baseline neural models, underscoring their ability to capture nuanced linguistic information.
Future work includes expanding the metric set to incorporate more discourse and pragmatic metrics, adapting PUCP-Metrix to other Spanish varieties, and integrating these metrics into pre-trained language models or NLP pipelines. Benchmarking on larger and more diverse tasks/datasets will further validate its robustness and support the development of specialized metric sets.
Limitations
The current evaluation has several limitations. Although PUCP-Metrix has been tested on multiple datasets, the experiments primarily focus on learner essays, children’s texts, and selected AuTexTification domains, leaving its performance on other genres and domains uncertain. Additionally, PUCP-Metrix depends heavily on spaCy-based linguistic processing and external lexicons (e.g., psycholinguistic norms), so parsing errors and coverage gaps in these resources can directly affect the reliability of the computed metrics.
Appendix A List of metrics in PUCP-Metrix
Category Metric Description Category Metric Description Descriptive Indices DESPC: Paragraph count Syntactic Complexity Indices SYNNP: Mean number of modifiers per noun phrase DESPCi: Paragraph count incidence per 1000 words SYNLE: Mean number of words before main verb DESSC: Sentence count SYNMEDwrd: Minimal edit distance of words between adjacent sentences DESSCi: Sentence count incidence per 1000 words SYNMEDlem: Minimal edit distance of lemmas between adjacent sentences DESWC: Word count (alphanumeric words) SYNMEDpos: Minimal edit distance of POS tags between adjacent sentences DESWCU: Unique word count SYNCLS1: Ratio of sentences with 1 clause DESWCUi: Unique word count incidence per 1000 words SYNCLS2: Ratio of sentences with 2 clauses DESPL: Average paragraph length (sentences per paragraph) SYNCLS3: Ratio of sentences with 3 clauses DESPLd: Standard deviation of paragraph length SYNCLS4: Ratio of sentences with 4 clauses DESSL: Average sentence length (words per sentence) SYNCLS5: Ratio of sentences with 5 clauses DESSLd: Standard deviation of sentence length SYNCLS6: Ratio of sentences with 6 clauses DESSNSL: Average sentence length excluding stopwords SYNCLS7: Ratio of sentences with 7 clauses DESSNSLd: Standard deviation of sentence length excluding stopwords Syntactic Pattern Density Indices DRNP: Noun phrase density per 1000 words DESSLmax: Maximum sentence length DRNPc: Noun phrase count DESSLmin: Minimum sentence length DRVP: Verb phrase density per 1000 words DESWLsy: Average syllables per word DRVPc: Verb phrase count DESWLsyd: Standard deviation of syllables per word DRNEG: Negation expression density per 1000 words DESCWLsy: Average syllables per content word DRNEGc: Negation expression count DESCWLsyd: Standard deviation of syllables per content word DRGER: Gerund form density per 1000 words DESCWLlt: Average letters per content word DRGERc: Gerund count DESCWLltd: Standard deviation of letters per content word DRINF: Infinitive form density per 1000 words DESWLlt: Average letters per word DRINFc: Infinitive count DESWLltd: Standard deviation of letters per word DRCCONJ: Coordinating conjunction density per 1000 words DESWNSLlt: Average letters per word (excluding stopwords) DRCCONJc: Coordinating conjunction count DESWNSLltd: Standard deviation of letters per word (excluding stopwords) DRSCONJ: Subordinating conjunction density per 1000 words DESLLlt: Average letters per lemma DRSCONJc: Subordinating conjunction count DESLLltd: Standard deviation of letters per lemma Connective Indices CNCAll: All connectives incidence per 1000 words Readability Indices RDFHGL: Fernández-Huertas Grade Level CNCCaus: Causal connectives incidence per 1000 words RDSPP: Szigriszt-Pazos Perspicuity CNCLogic: Logical connectives incidence per 1000 words RDMU: Readability µ index CNCADC: Adversative connectives incidence per 1000 words RDSMOG: SMOG index CNCTemp: Temporal connectives incidence per 1000 words RDFOG: Gunning Fog index CNCAdd: Additive connectives incidence per 1000 words RDHS: Honoré Statistic Word Information Indices WRDCONT: Content word incidence per 1000 words RDBR: Brunet index WRDCONTc: Content word count Referential Cohesion Indices CRFNO1: Noun overlap between adjacent sentences WRDNOUN: Noun incidence per 1000 words CRFAO1: Argument overlap between adjacent sentences WRDNOUNc: Noun count CRFSO1: Stem overlap between adjacent sentences WRDVERB: Verb incidence per 1000 words CRFCWO1: Content word overlap between adjacent sentences (mean) WRDVERBc: Verb count CRFCWO1d: Content word overlap between adjacent sentences (std dev) WRDADJ: Adjective incidence per 1000 words CRFANP1: Anaphore overlap between adjacent sentences WRDADJc: Adjective count CRFNOa: Noun overlap between all sentences WRDADV: Adverb incidence per 1000 words CRFAOa: Argument overlap between all sentences WRDADVc: Adverb count CRFSOa: Stem overlap between all sentences WRDPRO: Personal pronoun incidence per 1000 words CRFCWOa: Content word overlap between all sentences (mean) WRDPROc: Personal pronoun count CRFCWOad: Content word overlap between all sentences (std dev) WRDPRP1s: First person singular pronoun incidence per 1000 words CRFANPa: Anaphore overlap between all sentences WRDPRP1sc: First person singular pronoun count Lexical Diversity Indices LDTTRa: Type-token ratio for all words WRDPRP1p: First person plural pronoun incidence per 1000 words LDTTRcw: Type-token ratio for content words WRDPRP1pc: First person plural pronoun count LDTTRno: Type-token ratio for nouns WRDPRP2s: Second person singular pronoun incidence per 1000 words LDTTRvb: Type-token ratio for verbs WRDPRP2sc: Second person singular pronoun count LDTTRadv: Type-token ratio for adverbs WRDPRP2p: Second person plural pronoun incidence per 1000 words LDTTRadj: Type-token ratio for adjectives WRDPRP2pc: Second person plural pronoun count LDTTRLa: Type-token ratio for all lemmas WRDPRP3s: Third person singular pronoun incidence per 1000 words LDTTRLno: Type-token ratio for noun lemmas WRDPRP3sc: Third person singular pronoun count LDTTRLvb: Type-token ratio for verb lemmas WRDPRP3p: Third person plural pronoun incidence per 1000 words LDTTRLadv: Type-token ratio for adverb lemmas WRDPRP3pc: Third person plural pronoun count LDTTRLadj: Type-token ratio for adjective lemmas Psycholinguistic Indices PSYC: Overall concreteness ratio LDTTRLpron: Type-token ratio for pronouns PSYC0: Very low concreteness ratio (1-2.5) LDTTRLrpron: Type-token ratio for relative pronouns PSYC1: Low concreteness ratio (2.5-4) LDTTRLipron: Type-token ratio for indefinite pronouns PSYC2: Medium concreteness ratio (4-5.5) LDTTRLifn: Type-token ratio for functional words PSYC3: High concreteness ratio (5.5-7) LDMLTD: Measure of Textual Lexical Diversity (MTLD) PSYIM: Overall imageability ratio LDVOCd: Vocabulary Complexity Diversity (VoCD) PSYIM0: Very low imageability ratio (1-2.5) LDMaas: Maas index PSYIM1: Low imageability ratio (2.5-4) LDDno: Noun density PSYIM2: Medium imageability ratio (4-5.5) LDDvb: Verb density PSYIM3: High imageability ratio (5.5-7) LDDadv: Adverb density PSYFM: Overall familiarity ratio LDDadj: Adjective density PSYFM0: Very low familiarity ratio (1-2.5) Word Frequency Indices WFRCno: Rare noun count PSYFM1: Low familiarity ratio (2.5-4) WFRCnoi: Rare noun incidence per 1000 words PSYFM2: Medium familiarity ratio (4-5.5) WFRCvb: Rare verb count PSYFM3: High familiarity ratio (5.5-7) WFRCvbi: Rare verb incidence per 1000 words PSYAoA: Overall age of acquisition ratio WFRCadj: Rare adjective count PSYAoA0: Very early acquisition ratio (1-2.5) WFRCadji: Rare adjective incidence per 1000 words PSYAoA1: Early acquisition ratio (2.5-4) WFRCadv: Rare adverb count PSYAoA2: Medium acquisition ratio (4-5.5) WFRCadvi: Rare adverb incidence per 1000 words PSYAoA3: Late acquisition ratio (5.5-7) WFRCcw: Rare content word count PSYARO: Overall arousal ratio WFRCcwi: Rare content word incidence per 1000 words PSYARO0: Very low arousal ratio (1-3) WFRCcwd: Distinct rare content word count PSYARO1: Low arousal ratio (3-5) WFRCcwdi: Distinct rare content word incidence per 1000 words PSYARO2: Medium arousal ratio (5-7) WFMcw: Mean frequency of content words PSYARO3: High arousal ratio (7-9) WFMw: Mean frequency of all words PSYVAL: Overall valence ratio WFMrw: Mean frequency of rarest words per sentence PSYVAL0: Very negative valence ratio (1-4) WFMrcw: Mean frequency of rarest content words per sentence PSYVAL1: Negative valence ratio (3-5) Semantic Cohesion Indices SECLOSadj: LSA overlap between adjacent sentences (mean) PSYVAL2: Positive valence ratio (5-7) SECLOSadjd: LSA overlap between adjacent sentences (std dev) PSYVAL3: Very positive valence ratio (7-9) SECLOSall: LSA overlap between all sentences (mean) Textual Simplicity Indices SECLOSalld: LSA overlap between all sentences (std dev) TSSRsh: Ratio of short sentences (<11 words) SECLOPadj: LSA overlap between adjacent paragraphs (mean) TSSRmd: Ratio of medium sentences (11-12 words) SECLOPadjd: LSA overlap between adjacent paragraphs (std dev) TSSRlg: Ratio of long sentences (13-14 words) SECLOSgiv: LSA overlap between given and new sentences (mean) TSSRxl: Ratio of very long sentences ($\geq$ 15 words) SECLOSgivd: LSA overlap between given and new sentences (std dev)
Table 5: List of linguistic metrics implemented in PUCP-Metrix
Appendix B Datasets for Automated Readability Assessment and Machine-generated Text Detection
B.1 Automated Readability Assessment
  • CAES (Corpus de Aprendices del Español) [^4] Available at https://galvan.usc.es/caes/ Parodi (2015). This corpus consists of essays written by learners of Spanish as a foreign language. Each document is annotated with a CEFR level (A1–C2). Following Vásquez-Rodríguez et al. (2022), we map A1–B1 to "simple" and B2–C2 to "complex" for the 2-label schema, and A1-A2 to "basic", B1-B2 to "intermediate" and C1-C2 to "advanced" for the 3-label schema.
  • Coh-Metrix-Esp Quispesaravia et al. (2016). This dataset is a collection of short Spanish stories that includes children’s tales and texts intended for adults. It provides explicit simple and complex labels, directly aligned to our 2-label schema and to "basic" vs "advanced" in the 3-label schema.
  • Kwiziq [^5] The platform is available at https://www.kwiziq.com/. Kwiziq is an online language-learning platform that offers graded Spanish readings labeled with CEFR levels. We use the available data proposed by Vásquez-Rodríguez et al. (2022) and map the CEFR annotations to our 2- and 3-label classification schemes using the same criteria.
  • HablaCultura. This dataset comprises educational readings sourced from the HablaCultura platform [^6] Available at https://hablacultura.com/, where each text is labeled by instructors with CEFR levels. We use the same level mappings used by Vásquez-Rodríguez et al. (2022).
B.2 Machine-generated Text Detection
Human-generated texts in AuTexTification were sourced from publicly available datasets, including MultiEURLEX Chalkidis et al. (2021) (legal), XLSUM/MLSUM Hasan et al. (2021); Scialom et al. (2020) (news), COAR/COAH Molina-González et al. (2014) (reviews), XLM-Tweets Barbieri et al. (2022) and TSD Leis et al. (2019) (tweets), and WikiLingua Ladhak et al. (2020) (how-to articles). Machine-generated texts were produced using six large language models: three from the BLOOM family (BLOOM-1B7777Available at https://huggingface.co/bigscience/bloom-1b7., BLOOM-3B [^8] Available at https://huggingface.co/bigscience/bloom-3b., BLOOM-7B1999Available at https://huggingface.co/bigscience/bloom-7b1.) and three from the GPT-3 family (babbage, curie, text-davinci-003).
Appendix C Feature Analysis
We applied Anova over our dataset using all the metrics. We set a p-value of 0.05 and remove the features that do not make contribution for our analysis.
Figure 1 shows a heatmap representing the coverage of linguistic categories along the ranking, i.e., the distribution of linguistic features as more signals are included. Overall, the contribution of linguistic features varies across tasks. For machine-generated content detection, top-ranked signals are dominated by word frequency, readability, and semantic cohesion metrics. In contrast, descriptive and connective metrics play a more limited role and appear only at later ranks.
For ARA tasks, the importance shifts toward descriptive features, syntactic pattern density, readability, syntactic complexity, and textual simplicity metrics. Conversely, semantic cohesion and connective metrics are comparatively less important.
Figure 1: Category Coverage along the ranking for PUCP-Metrix